state the slope and y-intercept : 7y=4x-7
Found 2 solutions by ewatrrr, sachi:
Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
7y=4x-7 0r y = (4/7)x - 1, m = 4/7 and -1 is the y-intercept

Answer by sachi(548)

7y=4x-7
or y=4/7*x-1
so slope=4/7 & y intercept=-1
ans